OpenStudy (anonymous):

Express the given radian measure in degree measure to the nearest degree: 2

OpenStudy (phi):

same thing: multiply by 180/pi: \[ 2 \cdot \frac{180}{ \pi} = \frac{360}{\pi}\] you can use a calculator to find the final answer, and round to the nearest degree